Doubling time of an investment earning 2.5 interest if interest is compounded quarterly?
Answer by lwsshak3(11628)

Doubling time of an investment earning 2.5% interest if interest is compounded quarterly?
***
Compound interest formula: A=P(1+i)^n, P=initial investment, i=interest rate per compounding period, n=number of periods, A=amt after n-periods
For given problem:
A/P=(1+i)^n=2
i=.025/4=.00625
1+i=1.00625
..
1.00625^n=2
nlog(1.00625)=log(2)
n=log(2)/log(1.00625)
n≈111.25 periods
n≈28 yrs